# Optimization Proxies using Limited Labeled Data and Training Time – A Semi-Supervised Bayesian Neural Network Approach

### v1

@@ Abstract
Constrained optimization problems arise in various engineering system operations such as inventory management and electric power grids. However, the requirement to repeatedly solve such optimization problems with uncertain parameters poses a significant computational challenge. This work introduces a learning scheme using Bayesian Neural Networks (BNNs) to solve constrained optimization problems under limited labeled data and restricted model training times. We propose a semi-supervised BNN for this practical but complex regime, wherein training commences in a sandwiched fashion, alternating between a supervised learning step (using labeled data) for minimizing cost, and an unsupervised learning step (using unlabeled data) for enforcing constraint feasibility. Both supervised and unsupervised steps use a Bayesian approach, where Stochastic Variational Inference is employed for approximate Bayesian inference. We show that the proposed semi-supervised learning method outperforms conventional BNN and deep neural network (DNN) architectures on important non-convex constrained optimization problems from energy network operations, achieving up to a tenfold reduction in expected maximum equality gap and halving the optimality and inequality (feasibility) gaps, without requiring any correction or projection step. By leveraging the BNN’s ability to provide posterior samples at minimal computational cost, we demonstrate that a Selection via Posterior (SvP) scheme can further reduce equality gaps by more than 10%. We also provide tight and practically meaningful probabilistic confidence bounds that can be constructed using a low number of labeled testing data and readily adapted to other applications.
@@ 1 Introduction
Constrained optimization problems are fundamental in the optimal operation of various engineering systems, such as supply chains, transportation networks, and power grids. Learning a forward mapping between the inputs and outputs of these problems can significantly reduce computational burdens, especially when rapid solutions are required, such as in electricity market clearing or real-time transportation planning.
Recent advancements in machine learning (ML) have led to considerable efforts to solve optimization problems using deep neural networks (DNNs) Khadivi et al. (2023); Kotary et al. (2021); Fajemisin et al. (2024). The idea of learning input-to-output mappings has been explored via supervised and unsupervised methods, particularly in power system applications Zamzam & Baker (2020); Donti et al. (2021); Park & Van Hentenryck (2023); Fioretto et al. (2020); Kotary et al. (2021); Rolnick et al. (2022). Additionally, constraint penalization approaches have been proposed to enforce feasibility in predicted outputs within DNN loss functions AI4OPT (2023).
[…]
In practice, the input space of the problem in many applications is dynamic and changes constantly, necessitating frequent retraining and evaluation for ML models to remain effective amid distributional shifts in the input data. In the context of longer horizon planning problems involving bi-level optimization problems, these ML proxies serve as subroutines to simulate stochastic decision-making processes for given first-level decisions Ibrahim et al. (2020). Consequently, these models must be adaptable in both training and testing to changing problem inputs and parameters. Minimizing or limiting the time required for learning input/output mappings in optimization problems is thus crucial.
Estimating the generalization error over the testing dataset is another critical aspect, and is particularly relevant to planning and operations of engineered systems where the system must obey the physical and safety limits. As a result, when limited labeled data is available, one must rely on concentration bounds such as Hoeffding’s inequality Sridharan (2002); Hoeffding (1994) to develop expected error bounds using finite out-of-sample testing data. However, these bounds are often loose and impractical, creating the need for tighter expected error bounds with limited labeled testing data.
@@
Motivated by the preceding discussion, this paper considers the problem of designing optimization proxies with guaranteed confidence bounds in the setting of limited total time requirement and limited labeled sample availability. For practical applicability, the total time must account for training and testing data generation, model training time and prediction time. Our major contribution is the development of a semi-supervised Bayesian Neural Network (BNN) approach for this setting, that can be used to give tight confidence bounds on predictions. First, we propose using BNNs instead of DNNs for learning input-to-output mappings, as they provide intrinsic uncertainty quantification and allow the integration of prior beliefs Papamarkou et al. (2024). Second, we introduce a Sandwich learning method for BNN, which integrates unlabeled data into training through feasibility-based data augmentation. This approach enforces feasibility without requiring solved labeled instances. Third, we utilize the predictive variance information provided by BNNs to develop tight and practically useful expected error bounds using Bernstein concentration bounds Audibert et al. (2007). We intentionally restrict ourselves to 10 minutes of training time on a single CPU core to demonstrate the effectiveness of the proposed learning scheme under low-data, low-compute settings. For various power grid optimization problem instances, we show that (i) supervised BNNs outperform standard supervised DNN approaches under limited training time; (ii) the proposed Sandwich BNN enforces feasibility better than supervised BNNs without requiring additional computation time for training or data generation; and (iii) the Bernstein bound-based expected error bounds are tight and practically useful for constraint satisfaction studies without extra computational effort.
@@ 1.1 Related Work
In recent years, Deep Neural Networks (DNNs) have been applied to solve various optimization problems with physics-based constraints, particularly in energy networks Zamzam & Baker (2020); Gupta et al. (2022); Donti et al. (2021); Singh et al. (2021); Park & Van Hentenryck (2023); Kotary et al. (2021). The primary motivation is to replace time-consuming optimization algorithms with machine learning proxies, enabling instantaneous solutions to a large number of problem instances Park & Van Hentenryck (2023); Donti et al. (2021); Gupta et al. (2022); Zamzam & Baker (2020).
Outside the realm of optimization proxies, several semi-supervised learning methods have been proposed to leverage unlabeled data for improving ML model performance Yang et al. (2022). These approaches include augmenting unlabeled data with inexpensive pseudo-labels and developing unsupervised loss functions to be minimized alongside supervised loss functions Sharma et al. (2023); Yang et al. (2022). Data augmentation has been used in image classification with Bayesian Neural Networks (BNNs) using the notion of semantic similarity Sharma et al. (2023). However, this concept is not readily extensible to ML proxies for constrained optimization problems, where semantic similarity is hard to quantify for input variations leading to changes in the output. To address this challenge, we propose a feasibility-based data augmentation scheme that relates directly to the constraints of the optimization problem. To the best of our knowledge, these ideas have not been explored in the context of BNN algorithms for solving large-scale optimization problems. A related but distinct line of work involves loss function-based prior design for output constraint satisfaction Sam et al. (2024); Yang et al. (2020).
@@ 2.1 Problem Setup and Assumptions
We assume that for all $\forall\mathbf{x}\in\mathcal{X}$, i.e., for any input vector in the set $\mathcal{X}$ ($\mathcal{X}$ could be as simple as a hyper-rectangle), there exists at least one feasible solution to problem equation 1a. Let $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}^{\star}_{i})\}_{i=1}^{N}$ denote the labeled dataset, where $\mathbf{y}^{\star}_{i}$ is the optimal solution obtained by solving optimization problem equation 1a for each $\mathbf{x}_{i}$. Assuming that sampling the input vector $\mathbf{x}$ is inexpensive, we construct an unlabeled dataset $\mathcal{D}^{u}=\{\mathbf{x}_{j}\}_{j=1}^{M}$.
Our goal is to develop a BNN surrogate that provides an approximate optimal value of the decision variables $\widehat{\mathbf{y}}_{t}$ for a given test input vector $\mathbf{x}_{t}\in\mathcal{X}$. This work falls under the category of developing optimization proxies or surrogates, where the machine learning model serves as a direct forward mapping between the input and output variables of an optimization problem (see Park & Van Hentenryck (2023)).
The paper proposes a semi-supervised framework to solve this problem, wherein training alternates between a supervised step—using labeled data $\mathcal{D}$ to minimize prediction error—and an unsupervised step—using unlabeled data $\mathcal{D}^{u}$ to enforce the feasibility of constraints in equation 1b and equation 1c. Both steps are implemented using a Bayesian Neural Network (BNN).
@@ 2.2 Bayesian Neural Network
In the supervised training of the BNN, the goal is to compute the posterior distribution over the weights given the labeled data $\mathcal{D}\equiv(\mathbf{x},\mathbf{y})$. This posterior is expressed as
$p(w\mid\mathbf{x},\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{x},w)\,p(w)$
Here, $p(\mathbf{y}\mid\mathbf{x},w)$ is the likelihood of the labeled data given the weights, and $p(w)$ is the prior over the weights. The posterior distribution $p(w\mid\mathbf{x},\mathbf{y})$ encapsulates the uncertainty about the weights after observing the data. Due to computational challenges in calculating the normalization constant of the posterior, approximate methods such as stochastic variational inference (SVI) with the mean-field assumption are employed the posterior distribution estimation (see Jospin et al. (2022)).
@@ 3 Semi-supervised Learning: Sandwich BNN
We start by defining a suitable likelihood function for the unsupervised learning process. To that end, we augment the unlabeled data $\mathcal{D}^{u}$ using the necessary feasibility conditions that the vector $\mathbf{y}$ must satisfy to be a solution of equation 1a. We propose a function $\mathcal{F}(\mathbf{y},\mathbf{x})$ to measure the feasibility of a candidate solution $\mathbf{y}$ for a given input $\mathbf{x}$. This function consists of two terms: one measuring the equality gap and the other measuring the one-sided inequality gap or violations. The relative emphasis on each term is determined by the parameters $\lambda_{e}$ and $\lambda_{i}$, respectively, i.e.,
@@
For any given feasible solution $\mathbf{y}_{c}$ for the optimization problem in equation 1a [^2] Not necessarily optimal for equation 1a., we have $\mathcal{F}(\mathbf{y}_{c},\mathbf{x})=0$ for the given input $\mathbf{x}\in\mathcal{X}$. Furthermore, because of our assumption in Sec. 2.1 that the problem in equation 1a has at least one feasible solution, the minimum value $\mathcal{F}(\cdot,\mathbf{x})=0$ for any $\mathbf{x}\in\mathcal{X}$. Therefore, we can augment the unlabeled dataset $\mathcal{D}^{u}$ to create a labeled feasibility dataset, i.e., $\mathcal{D}^{f}=\{(\mathbf{x}_{j},\mathcal{F}(\cdot,\mathbf{x})=0)\}_{j=1}^{M}$. Since input sampling is inexpensive, constructing this feasibility dataset $\mathcal{D}^{f}$ incurs no additional computational cost. Similar to the supervised step in Sec. 2.2, we now define a Gaussian likelihood for the unsupervised training step, with $\sigma_{u}^{2}$ as the noise variance for unsupervised learning and $\mathbf{x}_{j}\in\mathcal{D}^{f}$, as
@@ 3.1 Selection via Posterior (SvP)
where $\mathbf{Y}_{\cdot j}$ is the $j$-th column of the PPM, and $g_{i}(\cdot,\cdot)$ represents the $i$-th equality constraint function. [^3] In a general setting of constrained optimization problems, there may be multiple equality and inequality constraints. The output prediction corresponding to the weight sample $W^{\star}$ will have the minimum equality gap, and we term this process Selection via Posterior (SvP). Note that the numerical operation in equation 3 can be performed in parallel and has minimal computational cost compared to analytical projection methods in Zamzam & Baker (2020) that focus on projecting the prediction onto equation 1c to satisfy the inequality constraints in the problem equation 1a. Note that it is an application motivated design choice to emphasize enforcement of the equality gap by using the SvP in equation 3. This can easily be adapted to account for inequality constraints without any significant computational overhead.
@@ 4 Probabilistic Confidence Bounds
Ideally, we would like to evaluate our model on a large number of samples since, as $M\rightarrow\infty$, the error bound $\varepsilon\rightarrow 0$. However, increasing $M$ leads to a higher requirement for labeled data, which defeats the purpose of training using low labeled data  [^4] Note that the total labeled data requirement is the sum of training and testing samples, i.e., $N+M$. To address this issue, confidence inequalities are commonly used to provide PCBs, with Hoeffding’s inequality (Hoeffding (1994)) being one of the most widely used bounds. As stated in Appendix D, Hoeffding’s inequality assumes that the error is bounded (i.e., $|e_{i}|\leqslant R$ for all $i$) and provides PCBs whose tightness is governed by $M$, with the relationship $\varepsilon=R\sqrt{\frac{\log\left({2}/{\delta}\right)}{2M}}.$ However, the Hoeffding bound can often be too loose to be practically relevant.
To improve upon this, we propose using Bernstein’s inequality (see Audibert et al. (2007)) as the concentration bound, which utilizes the total variance in error (TVE) information along with $M$, under the same bounded error assumption. The main challenge in using the Bernstein bound is obtaining the TVE without extensive out-of-sample testing. One possible solution is to use the empirical Bernstein bound as in Audibert et al. (2007); Mnih et al. (2008), which employs the empirical variance of the error $\widehat{\mathbb{V}}_{e}$, obtained from the same $M$ testing samples, and accounts for the error in variance estimation by modifying the theoretical Bernstein inequality. Mathematically, the PCB using the Empirical Bernstein inequality is
@@
as given in Audibert et al. (2007); Mnih et al. (2008). Since the term under the square root depends on the empirical TVE $\widehat{\mathbb{V}}_{e}$ rather than $R$, the empirical Bernstein bound becomes tighter more quickly with increasing $M$ if $\widehat{\mathbb{V}}_{e}\ll R$.
To further tighen the bound, we propose using the Theoretical Bernstein bound Sridharan (2002) (Theorem 3 in Appendix D) with the Mean Predictive Variance (MPV) as a proxy for the TVE $\mathbb{V}_{e}$. The MPV is the mean of the predictive variance of testing samples, i.e., $\text{MPV}=\frac{1}{M}\sum_{k=1}^{M}\mathbb{V}_{W}\left[\mathbf{Y}_{i\cdot}^{k}\right],$ where variance $\mathbb{V}_{W}\left[\mathbf{Y}_{i\cdot}^{k}\right]$ for the $k^{th}$ test sample using entries of the posterior prediction matrix across columns, generated using posterior weights, for $i^{th}$ output variable. In principle, the MPV captures the expected variance in the predictions due to the posterior distribution of BNN weights. We hypothesize that with a constant multiplier $\alpha>1$,
$\alpha\,\text{MPV}\geqslant\mathbb{V}_{e}=\mathbb{E}_{M}\left[\mathbb{V}_{W}[e]\right]+\mathbb{E}_{W}\left[\mathbb{V}_{M}[e]\right]\geqslant\mathbb{V}_{|e|},$ (5)
where $\mathbb{E}_{M}$ and $\mathbb{E}_{W}$ denote expectations with respect to $M$ testing samples and posterior weight samples, respectively, and $\mathbb{V}_{M}[e]$ and $\mathbb{V}_{W}[e]$ denote the variance of the error with respect to $M$ testing samples and posterior weight samples, respectively. The equality in equation 5 follows from the law of total variance Blitzstein & Hwang (2019). $\mathbb{V}_{|e|}$ represents the variance of the absolute value of the error, which is lower than the variance of the error.
Notice that the first term of the TVE, $\mathbb{E}_{M}\left[\mathbb{V}_{W}[e]\right]$, is independent of the labeled testing dataset because the true output $y$ is constant with respect to posterior weight samples; thus, $\mathbb{V}_{W}[e]=\mathbb{V}_{W}[y-y^{t}]=\mathbb{V}_{W}[y^{t}]$. Furthermore, from the definition of MPV, we have $\text{MPV}=\mathbb{E}_{M}\left[\mathbb{V}_{W}[e]\right]$. As an example, if $\mathbb{E}_{W}\left[\mathbb{V}_{M}[e]\right]\leqslant\text{MPV}$, our hypothesis in equation 5 holds with $\alpha=2$. Consequently, we can use $2\times\text{MPV}$ as an upper bound for $\mathbb{V}_{e}$ in the Theoretical Bernstein bound (see Theorem 3 in Appendix D), which gives the error bound
@@
which is better than the Empirical Bernstein bounds. The hypothesis in equation 5 and the corresponding constant $\alpha$ can be computed by using application specific information or performing a meta-study like in Section 5.
@@
In the next section, we perform a meta-study using different BNN models on different test cases to show the performance of our proposed learning architecture as well as demonstrate that hypothesis equation 5 indeed holds for the proposed optimization proxy learning problems.
5 Numerical Results and Discussion
We test the proposed method on the Alternating Current Optimal Power Flow (ACOPF) problem, essential for the economic operation of electrical power grids Molzahn et al. (2019). Efficient ACOPF proxies can mitigate climate change by enabling higher renewable energy integration, improving system efficiency by minimizing losses and emissions, and enhancing grid resiliency against extreme weather conditions Rolnick et al. (2022). ACOPF is a constrained optimization problem with nonlinear equality constraints and double-sided inequality bounds. It aims to find the most cost-effective generator set points while satisfying demand and adhering to physical and engineering constraints. The inputs are active and reactive power demands; outputs include generator settings, voltage magnitudes, and phase angles at each bus. We adopt the standard ACOPF formulation Babaeinejadsarookolaee et al. (2019); Park & Van Hentenryck (2023); Coffrin et al. (2018) and benchmark our method using the open-source OPFDataset from Torch_Geometric, which provides numerous solved ACOPF instances (see Lovett et al. (2024)).
In our results, ‘Gap%’ denotes the average relative optimality gap compared to the objective values in the labeled testing instances. ‘Max Eq.’ and ‘Mean Eq.’ represent the maximum and mean equality gaps over all equality constraints, while ‘Max Ineq.’ and ‘Mean Ineq.’ indicate the same for inequality gaps in per unit, all averaged over testing instances. We compare our proposed method with the following state-of-the-art baseline supervised learning models available in the literature using the same labeled dataset and training time constraints, utilizing AI4OPT’s ML4OPF package AI4OPT (2023) (network and hyper-parameter details are in Appendix B):
  • Naïve MAE and Naïve MSE (Supervised): Use $l_{1}$-norm and $l_{2}$-norm loss functions, respectively, to measure differences between predicted and actual optimal solutions Park & Van Hentenryck (2023), incorporating a bound repair layer with a sigmoid function. The bound repair layer ensures that inequality constraints are always satisfied.
  • MAE + Penalty, MSE + Penalty, and LD (Supervised): Add penalty terms for constraint violations to the naïve MAE or MSE loss functions Park & Van Hentenryck (2023). The Lagrangian Duality (LD) method applies the $l_{1}$-norm as outlined in Fioretto et al. (2020); Park & Van Hentenryck (2023), and also uses a bound repair layer with a sigmoid function.
We exclude self-supervised constrained optimization methods like Primal-Dual Learning (PDL) Park & Van Hentenryck (2023) and DC3 Donti et al. (2021) due to their significantly higher training times and computational demands, which violate the premise of this paper. For example, PDL requires over 125 minutes of training time for the ACOPF problem on a 118-node power network using a Tesla RTX6000 GPU. Methods that require solving alternating current power flow to recover solutions (e.g., Zamzam & Baker (2020)) are also excluded, as they result in high prediction times compared to DNN or BNN forward passes555see prediction time studies in Donti et al. (2021) which suggest 10 times higher prediction time with power flow based projections (0̃.080 sec. compared to 0.001 sec. for DNN and 0.003 sec. per testing instance for proposed BNNs)..

### v3

@@ Abstract
Constrained optimization problems arise in various engineering systems such as inventory management and power grids. Standard deep neural network (DNN) based machine learning proxies are ineffective in practical settings where labeled data is scarce and training times are limited. We propose a semi-supervised Bayesian Neural Networks (BNNs) based optimization proxy for this complex regime, wherein training commences in a sandwiched fashion, alternating between a supervised learning step for minimizing cost, and an unsupervised learning step for enforcing constraint feasibility. We show that the proposed semi-supervised BNN outperforms DNN architectures on important non-convex constrained optimization problems from energy network operations, achieving up to a tenfold reduction in expected maximum equality gap and halving the inequality gaps. Further, the BNN’s ability to provide posterior samples is leveraged to construct practically meaningful probabilistic confidence bounds on performance using a limited validation data, unlike prior methods. The implementation code for this study is available at: BNN-OPF.
Optimization Proxies
@@ 1 Introduction
Constrained optimization problems are fundamental in the optimal operation of various engineering systems, such as supply chains, transportation networks, and power grids. Learning a forward mapping between the inputs and outputs of these problems can significantly reduce computational burdens, especially when rapid solutions are required, such as in electricity markets or real-time transportation planning.
Recent advancements in machine learning (ML) have led to considerable efforts to solve optimization problems using deep neural networks (DNNs) (Khadivi et al., 2025; Kotary et al., 2021; Fajemisin et al., 2024). The idea of learning input-to-output mappings has been explored via supervised and unsupervised methods, particularly in power system applications (Zamzam & Baker, 2020; Donti et al., 2021; Park & Van Hentenryck, 2023; Fioretto et al., 2020; Kotary et al., 2021; Rolnick et al., 2022a; Piloto et al., 2024). Additionally, constraint penalization approaches have been proposed to enforce feasibility in predicted outputs within DNN loss functions (AI4OPT, 2023).
Supervised DNN models rely on labeled datasets obtained by solving numerous (${}10^{4}$ or more) instances of optimization problems. This data generation step poses a significant limitation due to prohibitive computational times required for moderate to large problem instances, particularly if the system topology and other parameters change over time. For example, (Park & Van Hentenryck, 2023) report that generating labeled data for a medium-sized power grid problem takes over three hours [^1] See Table 4 in (Park & Van Hentenryck, 2023).. Unsupervised methods aim to address the labeled data generation issue (Donti et al., 2021; Park & Van Hentenryck, 2023); however, they often have high training time requirements, particularly due to the use of projection/correction steps to ensure feasibility within the framework (Donti et al., 2021; Gupta et al., 2022; Zamzam & Baker, 2020). Moreover, unsupervised methods still require large number of labeled data to perform validation and provide confidence bounds on error with respect to true solution.
Thus, it is important to note that practical performance of ML based optimization proxies needs to be qualified under constraints of both : (a) Total Labeled Data (training $+$ validation), and (b) Training time. This is true in the context of bi-level optimization problems such as in power grid planning where ML proxies may serve as subroutines to simulate decision-making processes for given first-level decisions (Ibrahim et al., 2020). Here, ML models must be adaptable in both training and validation to changing problem inputs and parameters. Minimizing or limiting both the time and data required for learning input/output mappings in optimization problems is thus crucial.
Estimating the generalization error of ML models over the testing dataset is another aspect that is particularly relevant to engineered systems where the system must obey physical and safety limits. When limited labeled data is available for validation, one relies on concentration results such as Hoeffding’s inequality (Sridharan, 2002; Hoeffding, 1994) to develop error bounds using finite out-of-sample data. However, these bounds are often loose and impractical, and creates the need for frameworks that enable tighter expected error bounds with limited labeled validation data.
@@
Motivated by the preceding discussion, this paper considers the problem of designing optimization proxies with improved confidence bounds in the setting of limited training time requirement and limited labeled data availability. Our major contribution is the development of a Bayesian Neural Network (BNN) coupled with a semi-supervised training approach for this setting, that can be used to give tight confidence bounds on predictions. First, we propose the use of BNNs instead of DNNs for learning input-to-output mappings, as they provide intrinsic uncertainty quantification and allow the integration of prior beliefs (Papamarkou et al., 2024). Second, we introduce a Sandwich learning method for BNN, which integrates unlabeled data into training through feasibility-based data augmentation. This approach enforces feasibility without requiring more labeled instances. Third, we utilize the predictive variance information provided by BNNs to develop tight and practically useful expected error bounds using Bernstein concentration bounds (Audibert et al., 2007). We intentionally restrict ourselves to 1000 training instances and 10 minutes of training tim on a single CPU core to demonstrate the effectiveness of the proposed learning scheme under low-data, low-compute settings. For various power grid optimization problem instances (57-Bus, 118-Bus, 500-Bus and 2000-Bus), we show that (i) supervised BNNs outperform standard supervised DNN approaches under limited training time and data; (ii) the proposed Sandwich BNN enforces feasibility better than supervised BNNs without requiring additional training time or data; and (iii) the Bernstein bound-based expected error bounds are tight and practically useful for constraint satisfaction studies without extra computational effort. The proposed BNN-based approaches achieve at least an order of magnitude lower maximum equality gap compared to state-of-the-art DNN models, without compromising the optimality gap.
@@ 1.1 Related Work
In recent years, Deep Neural Networks (DNNs) have been applied to solve various optimization problems with physics-based constraints, particularly in energy networks (Zamzam & Baker, 2020; Gupta et al., 2022; Donti et al., 2021; Singh et al., 2021; Park & Van Hentenryck, 2023; Kotary et al., 2021). The primary motivation is to replace time-consuming optimization algorithms with machine learning proxies, enabling instantaneous solutions to a large number of problem instances (Park & Van Hentenryck, 2023; Donti et al., 2021; Gupta et al., 2022; Zamzam & Baker, 2020).
Outside the realm of optimization proxies, several semi-supervised learning methods have been proposed to leverage unlabeled data for improving ML model performance (Yang et al., 2022). These approaches include augmenting unlabeled data with inexpensive pseudo-labels and developing unsupervised loss functions to be minimized alongside supervised loss functions (Sharma et al., 2024; Yang et al., 2022). Data augmentation has been used in image classification with Bayesian Neural Networks (BNNs) using the notion of semantic similarity (Sharma et al., 2024). However, this concept is not readily extensible to ML proxies for constrained optimization problems, where semantic similarity is hard to quantify for input variations leading to changes in the output. To address this challenge, we propose a feasibility-based data augmentation scheme that relates directly to the constraints of the optimization problem. To the best of our knowledge, these ideas have not been explored in the context of BNN algorithms for solving large-scale optimization problems. A related but distinct line of work involves loss function-based prior design for output constraint satisfaction (Sam et al., 2024; Yang et al., 2020).
@@ 2.1 Problem Setup and Assumptions
We assume that for all $\forall\mathbf{x}\in\mathcal{X}$, i.e., for any input vector in the set $\mathcal{X}$ ($\mathcal{X}$ could be as simple as a hyper-rectangle), there exists at least one feasible solution to problem (1a). Let $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}^{\star}_{i})\}_{i=1}^{N}$ denote the labeled dataset, where $\mathbf{y}^{\star}_{i}$ is the optimal solution obtained by solving optimization problem (1a) for each $\mathbf{x}_{i}$. Assuming that sampling the input vector $\mathbf{x}$ is inexpensive, we construct an unlabeled dataset $\mathcal{D}^{u}=\{\mathbf{x}_{j}\}_{j=1}^{M}$.
Our goal is to develop a BNN surrogate that provides an approximate optimal value of the decision variables $\widehat{\mathbf{y}}_{t}$ for a given test input vector $\mathbf{x}_{t}\in\mathcal{X}$. This work falls under the category of developing optimization proxies or surrogates, where the machine learning model serves as a direct forward mapping between the input and output variables of an optimization problem (see (Park & Van Hentenryck, 2023)).
The paper proposes a semi-supervised framework to solve this problem, wherein training alternates between a supervised step—using labeled data $\mathcal{D}$ to minimize prediction error—and an unsupervised step—using unlabeled data $\mathcal{D}^{u}$ to enforce the feasibility of constraints in (1b) and (1c). Both steps are implemented using a Bayesian Neural Network (BNN).
@@ 2.2 Bayesian Neural Network
In the supervised training of the BNN, the goal is to compute the posterior distribution over the weights given the labeled data $\mathcal{D}\equiv(\mathbf{x},\mathbf{y})$. This posterior is expressed as $p(w\mid\mathbf{x},\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{x},w)\,p(w)$. Here, $p(\mathbf{y}\mid\mathbf{x},w)$ is the likelihood of the labeled data given the weights, and $p(w)$ is the prior over the weights. The posterior distribution $p(w\mid\mathbf{x},\mathbf{y})$ encapsulates the uncertainty about the weights after observing the data. Due to computational challenges in calculating the normalization constant of the posterior, approximate methods such as stochastic variational inference (SVI) with the mean-field assumption are employed the posterior distribution estimation (see (Jospin et al., 2022)).
@@ 3 Semi-supervised Learning: Sandwich BNN
We start by defining a suitable likelihood function for the unsupervised learning process. To that end, we augment the unlabeled data $\mathcal{D}^{u}$ using the necessary feasibility conditions that the vector $\mathbf{y}$ must satisfy to be a solution of (1a). We propose a function $\mathcal{F}(\mathbf{y},\mathbf{x})$ to measure the feasibility of a candidate solution $\mathbf{y}$ for a given input $\mathbf{x}$. This function consists of two terms: one measuring the equality gap and the other measuring the one-sided inequality gap or violations. The relative emphasis on each term is determined by the parameters $\lambda_{e}$ and $\lambda_{i}$, respectively, i.e.,
@@
For any given feasible solution $\mathbf{y}_{c}$ for the optimization problem in (1a) [^2] Not necessarily optimal for (1a)., we have $\mathcal{F}(\mathbf{y}_{c},\mathbf{x})=0$ for the given input $\mathbf{x}\in\mathcal{X}$. Furthermore, because of our assumption in Sec. 2.1 that the problem in (1a) has at least one feasible solution, the minimum value $\mathcal{F}(\cdot,\mathbf{x})=0$ for any $\mathbf{x}\in\mathcal{X}$. Therefore, we can augment the unlabeled dataset $\mathcal{D}^{u}$ to create a labeled feasibility dataset, i.e., $\mathcal{D}^{f}=\{(\mathbf{x}_{j},\mathcal{F}(\cdot,\mathbf{x})=0)\}_{j=1}^{M}$. Since input sampling is inexpensive, constructing this feasibility dataset $\mathcal{D}^{f}$ incurs no additional computational cost. Similar to the supervised step in Sec. 2.2, we now define a Gaussian likelihood for the unsupervised training step, with $\sigma_{u}^{2}$ as the noise variance for unsupervised learning and $\mathbf{x}_{j}\in\mathcal{D}^{f}$, as
@@ 3.1 Selection via Posterior (SvP)
where $\mathbf{Y}_{\cdot j}$ is the $j$-th column of the PPM, and $g_{i}(\cdot,\cdot)$ represents the $i$-th equality constraint function. [^3] In a general setting of constrained optimization problems, there may be multiple equality and inequality constraints. The output prediction corresponding to the weight sample $W^{\star}$ will have the minimum equality gap, and we term this process Selection via Posterior (SvP). Note that the numerical operation in (3) can be performed in parallel and has minimal computational cost compared to analytical projection methods in (Zamzam & Baker, 2020) that focus on projecting the prediction onto (1c) to satisfy the inequality constraints in the problem (1a). Note that it is an application motivated design choice to emphasize enforcement of the equality gap by using the SvP in (3). This can easily be adapted to account for inequality constraints without any significant computational overhead.
@@ 4 Probabilistic Confidence Bounds
Ideally, we would like to evaluate our model on a large number of samples since, as $M\rightarrow\infty$, the error bound $\varepsilon\rightarrow 0$. However, increasing $M$ leads to a higher requirement for labeled data, which defeats the purpose of training using low labeled data  [^4] Note that the total labeled data requirement is the sum of training and testing samples, i.e., $N+M$. To address this issue, confidence inequalities are commonly used to provide PCBs, with Hoeffding’s inequality ((Hoeffding, 1994)) being one of the most widely used bounds. As stated in Appendix E, Hoeffding’s inequality assumes that the error is bounded (i.e., $|e_{i}|\leqslant R$ for all $i$) and provides PCBs whose tightness is governed by $M$, with the relationship $\varepsilon=R\sqrt{\frac{\log\left({2}/{\delta}\right)}{2M}}.$ However, the Hoeffding’s bound can often be too loose to be practically relevant.
To improve upon this, we propose using Bernstein’s inequality (see (Audibert et al., 2007)) as the concentration bound, which utilizes the total variance in error (TVE) information along with $M$, under the same bounded error assumption. The main challenge in using the Bernstein bound is obtaining the TVE without extensive out-of-sample testing. One possible solution is to use the empirical Bernstein bound as in (Audibert et al., 2007; Mnih et al., 2008), which employs the empirical variance of the error $\widehat{\mathbb{V}}_{e}$, obtained from the same $M$ testing samples, and accounts for the error in variance estimation by modifying the theoretical Bernstein inequality. Mathematically, the PCB using the Empirical Bernstein inequality is
@@
as given in (Audibert et al., 2007; Mnih et al., 2008). Since the term under the square root depends on the empirical TVE $\widehat{\mathbb{V}}_{e}$ rather than $R$, the empirical Bernstein bound becomes tighter more quickly with increasing $M$ if $\widehat{\mathbb{V}}_{e}\ll R$.
To further tighen the bound, we propose using the Theoretical Bernstein bound (Sridharan, 2002) (Theorem E.3 in Appendix E) with the Mean Predictive Variance (MPV) as a proxy for the TVE $\mathbb{V}_{e}$. The MPV is the mean of the predictive variance of testing samples, i.e., $\text{MPV}=\frac{1}{M}\sum_{k=1}^{M}\mathbb{V}_{W}\left[\mathbf{Y}_{i\cdot}^{k}\right],$ where variance $\mathbb{V}_{W}\left[\mathbf{Y}_{i\cdot}^{k}\right]$ for the $k^{th}$ test sample using entries of the posterior prediction matrix across columns, generated using posterior weights, for $i^{th}$ output variable. In principle, the MPV captures the expected variance in the predictions due to the posterior distribution of BNN weights. We hypothesize that with a constant multiplier $\alpha>1$,
$\alpha\,\text{MPV}\geqslant\mathbb{V}_{e}=\mathbb{E}_{M}\left[\mathbb{V}_{W}[e]\right]+\mathbb{V}_{W}\left[\mathbb{E}_{M}[e]\right]\geqslant\mathbb{V}_{|e|},$ (5)
where $\mathbb{E}_{M}$ and $\mathbb{E}_{W}$ denote expectations with respect to $M$ testing samples and posterior weight samples, respectively, and $\mathbb{V}_{M}[e]$ and $\mathbb{V}_{W}[e]$ denote the variance of the error with respect to $M$ testing samples and posterior weight samples, respectively. The equality in (5) follows from the law of total variance (Blitzstein & Hwang, 2019). $\mathbb{V}_{|e|}$ represents the variance of the absolute value of the error, which is lower than the variance of the error.
Notice that the first term of the TVE, $\mathbb{E}_{M}\left[\mathbb{V}_{W}[e]\right]$, is independent of the labeled testing dataset because the true output $y$ is constant with respect to posterior weight samples; thus, $\mathbb{V}_{W}[e]=\mathbb{V}_{W}[y-y^{t}]=\mathbb{V}_{W}[y^{t}]$. Furthermore, from the definition of MPV, we have $\text{MPV}=\mathbb{E}_{M}\left[\mathbb{V}_{W}[e]\right]$. As an example, if $\mathbb{V}_{W}\left[\mathbb{E}_{M}[e]\right]\leqslant\text{MPV}$, our hypothesis in (5) holds with $\alpha=2$. Consequently, we can use $2\times\text{MPV}$ as an upper bound for $\mathbb{V}_{e}$ in the Theoretical Bernstein bound (see Theorem E.3 in Appendix E), which gives the error bound
@@
which is better than the Empirical Bernstein bounds. The hypothesis in (5) and the corresponding constant $\alpha$ can be computed by using application specific information or performing a meta-study like in Section 5.
@@
In the next section, we perform a meta-study using different BNN models on different test cases to show the performance of our proposed learning architecture as well as demonstrate that hypothesis (5) indeed holds for the proposed optimization proxy learning problems.